OpenStudy (anonymous):

Integrate x^2+y^2+z^2 over the cylinder x^2+y^2 ≤2, 2≤z≤3. are the bounds of the cylinder the bounds you integrate by?

OpenStudy (turingtest):

yes, in cylindrical coordinates (x,y,z,) becomes (r, theta, z) r goes from 0 to 2 theta goes from 0 to 2pi z goes from 2 to 3 so those are your bounds

OpenStudy (anonymous):

ok thanks just wanted to make sure i was on the right track. thanks again!

OpenStudy (anonymous):

so would my integration look like this? \[\int\limits_{2}^{3}\int\limits_{0}^{2}\int\limits_{0}^{2\Pi} r+z^2 dzdrd \Theta\]